Consider a particle moving in a central force field 


f(u) = - ku3


using standard notations u = 1/r. The quantity


d2u/dθ2 + u


(a) is proportional to u2


(b) is proportional to u


(c) is a constant


(d) is proportional to 1/u


(e) none of these


1
Expert's answer
2021-03-09T07:47:24-0500

Answer

According to centrel force motion

is given

"F(\frac{1}{u}) =-\frac{J^2u^2}{m}(\frac{d^2u}{d\theta^2}+u)"

Therefore when we get

"(\frac{d^2u}{d\theta^2}+u)=-F(\frac{1}{u})\frac{J^2u^2}{m}"

Therefore we can can say the quantity

"(\frac{d^2u}{d\theta^2}+u)" is directly inversely proportional to u2. so option e is correct.